decide whether the points are vertices of a right triangle..(1,2), (3,0), (3,3)
Answer by fractalier(6550)

To do this you need to find the slopes of the lines connecting these points.
Call the points #1, #2, and #3.
The slope of #1 & #2 is (0-2)/(3-1) = -1
The slope of #2 & #3 is (3-0)/(3-3) = infinity ==> vertical line
The slope of #3 & #1 is (2-3)/(1-3) = 1/2
None of these slopes is the negative reciprocal of another.
Therefore this is not a right triangle.